# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v4

@@ Appendix B Proof of Lemma 2
For the case of $\sqrt{(N_{x}\Delta)^{2}+(N_{y}\Delta)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$, i.e., $\epsilon_{n}=0,\;\forall n$, the SNR expression in (24) reduces to
@@
=$\overset{(d)}{=}$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\left.\frac{1}{\sqrt{1+y^{2}}}\operatorname{arctan}\left(\frac{x}{\sqrt{1+y^{2}}}\right)\mathrm{d}y\right|_{-\frac{N_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}$
@@
$\overset{(e)}{\approx}$ $\frac{1}{4}N_{x}\pi\delta r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\frac{1}{1+y^{2}}\mathrm{d}y$
$=$ $\left.\frac{1}{4}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(y\right)\right|_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}$
$=$ $\frac{1}{2}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(\frac{N_{y}\delta}{2}\right)$
$\overset{(f)}{\approx}$ $\frac{N_{x}N_{y}\pi^{2}\delta^{2}r^{2}}{16},$ (59)
where $(d)$ follows the integral formula 2.172 in [47], $(e)$ and $(f)$ hold by exploiting the linear approximation of $\operatorname{arctan}(x)\approx\frac{\pi}{4}x$, $-1\leq x\leq 1$ [42].
Thus, by substituting (B) into (58) and considering the conditions that $-1\leq\frac{N_{x}\delta}{2}\leq 1$ and $-1\leq\frac{N_{y}\delta}{2}\leq 1$, i.e., $N_{x},N_{y}\leq\frac{2}{\delta}$ or $\theta_{\mathrm{max}}\leq\frac{\pi}{4}$, Lemma 2 can be obtained.
@@ Appendix C Proof of Theorem 2
$\quad\quad\quad\quad\left.\int_{0}^{2\pi}\mathrm{d}\zeta\int_{D}^{R}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(\frac{l}{r}\right)-\theta_{\mathrm{max}}\right)}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l\right),$ (61)
@@
$\mathbb{D}_{2}=$ $2\pi r^{2}\int_{\frac{D}{r}}^{\frac{R}{r}}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(l\right)-\theta_{\mathrm{max}}\right)}{1+l^{2}}l\mathrm{d}l$
$=$ $2\pi r^{2}\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon.$ (63)
@@ Appendix D Proof of Lemma 3
$\mathbb{D}=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\operatorname{cos}(\epsilon-\theta_{\mathrm{max}})\operatorname{tan}\epsilon\mathrm{d}\epsilon$
$=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\left(\operatorname{cos}\theta_{\mathrm{max}}\operatorname{sin}\epsilon+\operatorname{sin}\theta_{\mathrm{max}}\frac{\operatorname{sin}^{2}\epsilon}{\operatorname{cos}\epsilon}\right)\mathrm{d}\epsilon$
$=$ $\operatorname{cos}\theta_{\mathrm{max}}\left(\operatorname{cos}\left(\operatorname{arctan}\left(\frac{D}{r}\right)\right)-\operatorname{cos}\left(\operatorname{arctan}\left(\frac{R}{r}\right)\right)\right)+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\int_{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{D}{r}\right)\right)}^{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{R}{r}\right)\right)}\frac{\epsilon^{2}}{1-\epsilon^{2}}\mathrm{d}\epsilon$
=$\overset{(g)}{=}$ $\operatorname{cos}\theta_{\mathrm{max}}\left(\operatorname{cos}\theta_{\mathrm{max}}-\operatorname{cos}\theta_{R}\right)+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\left.\left(\frac{1}{2}\operatorname{ln}\left(\frac{1+\epsilon}{1-\epsilon}\right)-\epsilon\right)\right|_{\operatorname{sin}\theta_{\mathrm{max}}}^{\operatorname{sin}\theta_{R}}$
@@
where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$, $(g)$ follows the integral formulas 2.147.1 and 2.111.6 in [47], and it has $\theta_{\mathrm{max}}=\operatorname{arctan}\left(\frac{D}{r}\right)$ when $R\geq D$.
Thus, by substituting (D) into (III-B), Lemma 3 can be obtained.

### v5

@@ Appendix B Proof of Lemma 2
For the case of $\sqrt{(N_{x}\Delta)^{2}+(N_{y}\Delta)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$, i.e., $\varpi_{n}=0,\;\forall n$, the SNR expression in (24) reduces to
@@
$\overset{(d)}{=}$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\left.\frac{1}{\sqrt{1+y^{2}}}\operatorname{arctan}\left(\frac{x}{\sqrt{1+y^{2}}}\right)\mathrm{d}y\right|_{-\frac{N_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}$
@@
$\overset{(e)}{\approx}$ $\frac{1}{4}N_{x}\pi\delta r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\frac{1}{1+y^{2}}\mathrm{d}y=\left.\frac{1}{4}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(y\right)\right|_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}$
$=$ $\frac{1}{2}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(\frac{N_{y}\delta}{2}\right)\overset{(f)}{\approx}\frac{N_{x}N_{y}\pi^{2}\delta^{2}r^{2}}{16},$ (59)
where $(d)$ follows the integral formula 2.172 in [10], $(e)$ and $(f)$ hold by exploiting the linear approximation of $\operatorname{arctan}(x)\approx\frac{\pi}{4}x$, $1\leq x\leq 1$ [23]. Thus, by substituting (B) into (58) and considering the conditions that $1\leq\frac{N_{x}\delta}{2}\leq 1$ and $1\leq\frac{N_{y}\delta}{2}\leq 1$, i.e., $N_{x},N_{y}\leq\frac{2}{\delta}$ or $\theta_{\mathrm{max}}\leq\frac{\pi}{4}$, Lemma 2 can be obtained.
@@ Appendix C Proof of Theorem 2
$\quad\quad\quad\left.\int_{0}^{2\pi}\mathrm{d}\zeta\int_{D}^{R}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(\frac{l}{r}\right)-\theta_{\mathrm{max}}\right)}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l\right),\hskip-5.69046pt$ (61)
@@
$\hskip-7.11317pt\mathbb{D}_{2}=$ $2\pi r^{2}\int_{\frac{D}{r}}^{\frac{R}{r}}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(l\right)-\theta_{\mathrm{max}}\right)}{1+l^{2}}l\mathrm{d}l$
$\hskip-7.11317pt=$ $2\pi r^{2}\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\varpi\operatorname{tan}(\varpi+\theta_{\mathrm{max}})\mathrm{d}\varpi.\hskip-5.69046pt$ (63)
@@ Appendix D Proof of Lemma 3
$\mathbb{D}=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\operatorname{cos}(\varpi-\theta_{\mathrm{max}})\operatorname{tan}\varpi\mathrm{d}\varpi$
$=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\left(\operatorname{cos}\theta_{\mathrm{max}}\operatorname{sin}\varpi+\operatorname{sin}\theta_{\mathrm{max}}\frac{\operatorname{sin}^{2}\varpi}{\operatorname{cos}\varpi}\right)\mathrm{d}\varpi$
$=$ $\operatorname{cos}\theta_{\mathrm{max}}\hskip-1.42271pt\left(\operatorname{cos}\left(\operatorname{arctan}\left(\hskip-1.42271pt\frac{D}{r}\right)\hskip-1.42271pt\right)\hskip-1.42271pt-\hskip-1.42271pt\operatorname{cos}\left(\operatorname{arctan}\left(\hskip-1.42271pt\frac{R}{r}\hskip-0.56917pt\right)\right)\right)\hskip-1.42271pt+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\int_{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{D}{r}\right)\right)}^{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{R}{r}\right)\right)}\frac{\varpi^{2}}{1-\varpi^{2}}\mathrm{d}\varpi$
$\overset{(g)}{=}$ $\operatorname{cos}\theta_{\mathrm{max}}\left(\operatorname{cos}\theta_{\mathrm{max}}-\operatorname{cos}\theta_{R}\right)+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\left.\left(\frac{1}{2}\operatorname{ln}\left(\frac{1+\varpi}{1-\varpi}\right)-\varpi\right)\right|_{\operatorname{sin}\theta_{\mathrm{max}}}^{\operatorname{sin}\theta_{R}}$
@@
where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$, $(g)$ follows the integral formulas 2.147.1 and 2.111.6 in [10], and it has $\theta_{\mathrm{max}}=\operatorname{arctan}\left(\frac{D}{r}\right)$ when $R\geq D$. Thus, by substituting (D) into (III-B), Lemma 3 can be obtained.
Appendix E Polarization-Aware Wideband Channel Model
Polarization is a fundamental characteristic of electromagnetic (EM) waves, and the polarization direction denotes the oscillation orientation of the electric field [3]. In RA systems, antenna rotation will change the polarization axis, which may lead to misalignment between transmit and receive polarizations and thereby induce polarization mismatch loss. To incorporate this effect into the RA channel model, we assume that both the BS and users employ linearly polarized antennas. As illustrated in Fig. 14, the electric field vector of an EM wave is always perpendicular to its propagation direction and lies in the plane formed by the antenna polarization axis (i.e., main polarization direction) and the propagation direction.
Figure 14: Illustration of polarization direction and electric field vector.
Assume that the main polarization direction of each RA is initially aligned with the $\tilde{x}-$axis in its local coordinate system $\tilde{o}-$$\tilde{x}\tilde{y}\tilde{z}$ as shown in Fig. 14, consistent with the pointing vector defined in (3). After antenna rotation, the polarization direction of RA $n$ can be expressed as
$\vec{\mathbf{p}}_{n}=[\operatorname{cos}\theta_{\mathrm{z},n}\operatorname{cos}\theta_{\mathrm{a},n},\operatorname{cos}\theta_{\mathrm{z},n}\operatorname{sin}\theta_{\mathrm{a},n},-\operatorname{sin}\theta_{\mathrm{z},n}]^{T}.$ (65)
The projection of $\vec{\mathbf{p}}_{n}$ onto the tangent plane orthogonal to the EM wave propagation from user $k$ to RA $n$ is given by
$\mathbf{p}_{\mathrm{U},k,n}^{\mathrm{r}}=\vec{\mathbf{p}}_{n}-(\vec{\mathbf{p}}_{n}^{T}\vec{\mathbf{u}}_{k,n})\vec{\mathbf{u}}_{k,n}.$ (66)
Similarly, the projection of $\vec{\mathbf{p}}_{n}$ onto the tangent plane orthogonal to the EM wave propagation from scatterer cluster $q$ to RA $n$ can be obtained and is denoted by $\mathbf{p}_{\mathrm{C},q,n}^{\mathrm{r}}$.
Let $\mathbf{p}_{\mathrm{U},k,n}^{\mathrm{t}}$ and $\mathbf{p}_{\mathrm{C},q,n}^{\mathrm{t}}$ denote the electric field vectors of the EM waves transmitted from user $k$ and reflected from scatterer cluster $q$, respectively. The channel gains accounting for polarization effects are then given by
$\varrho_{\mathrm{U},k,n}=(\mathbf{p}_{\mathrm{U},k,n}^{\mathrm{t}})^{T}\mathbf{p}_{\mathrm{U},k,n}^{\mathrm{r}},$ (67)
$\varrho_{\mathrm{C},q,n}=(\mathbf{p}_{\mathrm{C},q,n}^{\mathrm{t}})^{T}\mathbf{p}_{\mathrm{C},q,n}^{\mathrm{r}},$ (68)
for the user $k$-RA $n$ and cluster $q$-RA $n$ links, respectively.
By incorporating polarization effects, the channel gains of the LoS and NLoS links from user $k$ to RA $n$ are updated as
$\hskip-5.69046pt\tilde{g}_{k,0}(\vec{\mathbf{f}}_{n})$ $=\varrho_{\mathrm{U},k,n}\sqrt{g_{\mathrm{U},k}(\vec{\mathbf{f}}_{n})},$ (69)
$\hskip-5.69046pt{\tilde{g}_{k,q}}({\vec{\bf{f}}_{n}})$ $=\frac{{\varrho_{\mathrm{C},q,n}}}{{{r_{k,q}^{\mathrm{U-C}}}}}\sqrt{{\varsigma_{q}}\;{g_{{\rm{C}},q}}({\vec{\bf{f}}_{n}})}\;e^{j\chi_{q}},\;q=1,2,\ldots,Q.\hskip-8.5359pt$ (70)
Let ${\tau_{n,k,0}}=\frac{{{r_{k,n}^{\mathrm{B-U}}}}}{c}$ and ${\tau_{n,k,q}}=\frac{{{r_{q,n}^{\mathrm{B-C}}}+{r_{k,q}^{\mathrm{U-C}}}}}{c}$ denote the propagation delay of the “user $k$-RA nn” and “user $k$-cluster $q$-RA nn” links, respectively, where $c$ is the velocity of light. Before receiver filtering and sampling, the continuous-time baseband equivalent channel impulse response between user $k$ and RA $n$ is given by
${\tilde{h}_{k}}({{{\vec{\bf{f}}}_{n}},t})=\sum\limits_{q=0}^{Q}{{{\tilde{g}}_{k,q}}({{{\vec{\bf{f}}}_{n}}})}\;{e^{-j2\pi{f_{c}}{\tau_{n,k,q}}}}\delta\left({t-{\tau_{n,k,q}}}\right),$ (71)
where $f_{c}$ is the carrier frequency. The continuous-time Fourier transformation of ${\tilde{h}_{k}}({{{\vec{\bf{f}}}_{n}},t})$ is then obtained as [6]
${\tilde{H}_{k}}({{{\vec{\bf{f}}}_{n}},f})=\sum\limits_{q=0}^{Q}{{{\tilde{g}}_{k,q}}({{{\vec{\bf{f}}}_{n}}})}\;{e^{-j2\pi{f_{c}}\left({1+\frac{f}{{{f_{c}}}}}\right){\tau_{n,k,q}}}},$ (72)
which denotes the spatial-frequency channel response between user $k$ and RA $n$. The discrete-time frequency-domain channel used for OFDM demodulation can be obtained from (72) under Nyquist pulse shaping and with a sufficiently long CP. Let $\Delta f$ denote the OFDM subcarrier spacing, and $B=L\Delta f$ be the total bandwidth with $L$ subcarriers. The frequency of the $l$-th subcarrier is ${f_{l}}=(l-1)\Delta f$ with $l=1,2,\ldots,L$ [6]. Therefore, spatial-frequency channel response between user $k$ and the BS at subcarrier $l$ is given by
$\hskip-7.11317pt\tilde{\bm{h}}_{k,l}(\mathbf{F})\hskip-2.84544pt=\hskip-2.84544pt\left[\tilde{H}_{k}(\vec{\mathbf{f}}_{1},\hskip-1.42271ptf_{l}),\tilde{H}_{k}(\vec{\mathbf{f}}_{2},\hskip-1.42271ptf_{l}),\ldots,\tilde{H}_{k}(\vec{\mathbf{f}}_{N},\hskip-1.42271ptf_{l})\right]^{T}\hskip-5.69046pt\in\hskip-2.84544pt\mathbb{C}^{N\hskip-1.42271pt\times\hskip-1.42271pt1}.\hskip-5.69046pt$ (73)
To avoid inter-user interference, we assume each subcarrier is allocated to at most one user. Define a binary variable $a_{k,l}$}, which indicates whether subcarrier $l$ is allocated to user $k$, i.e., ${a_{k,l}}=1$ if subcarrier $l$ is assigned to user $k$, and ${a_{k,l}}=0$ otherwise. The subcarrier allocation constraints are given by
$\sum\limits_{k=1}^{K}{{a_{k,l}}}\leq 1,\;\forall l,$ (74)
and the transmit power constraint for user $k$ is
$\sum\limits_{l=1}^{L}{{a_{k,l}}}{P_{k,l}}\leq{P_{k}},\;\forall k,$ (75)
where $P_{k,l}$ denotes the transmit power of user $k$ at subcarrier $l$.
If ${a_{k,l}}=1$ and the BS has perfect synchronization (including symbol timing, frame alignment, and carrier frequency offset (CFO) compensation), the signal transmitted from user $k$ and received at the BS over subcarrier $l$ can be expressed as
${\tilde{\bf{y}}_{k,l}}={\tilde{\bf{h}}_{k,l}}\left({\bf{F}}\right)\sqrt{{P_{k,l}}}{s_{k,l}}+{{\bf{n}}_{l}},$ (76)
where $s_{k,l}$ denotes the information-bearing signal transmitted by user $k$ at subcarrier $l$, and $\mathbf{n}_{l}$ is the AWGN vector with variance $\tilde{\sigma}_{l}^{2}$. Since there is no inter-user interference over each subcarrier, the BS can apply the MRC beamforming, i.e., ${\tilde{\bf{v}}_{k,l}}=\frac{{{{\tilde{\bf{h}}}_{k,l}}\left({\bf{F}}\right)}}{{\|{{{\tilde{\bf{h}}}_{k,l}}\left({\bf{F}}\right)}\|}}$, to maximize the receive SNR. Thus, the achievable rate in bits per second per Hertz (bps/Hz) of user $k$ is given by
${\tilde{R}_{k}}=\frac{1}{{L+{L_{{\rm{CP}}}}}}\sum\limits_{l=1}^{L}{{a_{k,l}}{{\log}_{2}}\left({1+\frac{{{P_{k,l}}{{\|{{{\tilde{\bf{h}}}_{k,l}}\left({\bf{F}}\right)}\|}^{2}}}}{{\tilde{\sigma}_{l}^{2}}}}\right)},$ (77)
where $L_{{\rm{CP}}}$ is the length of the CP.
We formulate the following optimization problem to maximize the sum rate across all users and subcarriers by jointly optimizing the RA pointing matrix $\mathbf{F}$ and the subcarrier allocation $\{a_{k,l}\}$.
$\text{(P11):}\quad\max_{\mathbf{F},\{a_{k,l}\}}\quad$ $\sum_{k=1}^{K}\tilde{R}_{k}$ (78a)
s.t. $a_{k,l}\in\{0,1\},\;\forall k,l,$ (78b)
[…]
To solve the above problem, we can alternately optimize the pointing matrix $\mathbf{F}$ and the subcarrier allocation $\{a_{k,l}\}$ in an iterative manner. For a given $\mathbf{F}$, the subproblem of optimizing $\{a_{k,l}\}$ is a typical binary integer programming problem, which can be effectively solved using the CVX solver [5]. However, for a given $\{a_{k,l}\}$, the subproblem of optimizing $\mathbf{F}$ remains highly non-convex. In addition, the objective function (78a) is non-concave, and constraint (14c) is non-convex. To overcome these challenges, we first linearly approximate equality constraint (14c) to approximate the feasible region as a convex set. Then, the conditional gradient method is employed to obtain the solution of $\mathbf{F}$ by using a linear programming solver, such as linprog [24]. Based on the principle of the BCD method, the optimized $\mathbf{F}$ and $\{a_{k,l}\}$ can be obtained by alternatively solving the above two subproblems until convergence is attained.